Grade 12Algebra

Solve:
(3/5) + (2/7)
Solve:
(3/5) + (2/7)
Solve:
(3/5) + (2/7)

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

For addition of two unlike fractions, first change them to the like fractions.
LCM of 5, 7 = 35
Now, let us change each of the given fraction into an equivalent fraction having 35 as the denominator.
= [(3/5) × (7/7)] = (21/35)
= [(2/7) × (5/5)] = (10/35)
Now,
= (21/35) + (10/35)
= [(21 + 10)/35]
= (31/35)

Thanks